OpenStudy (anonymous):

cos theta-tan theta cos theta=0 a. 0, pi/4, pi, 5pi/4 b.pi/4, 5pi/4 c. pi/2, 3pi/4, 3pi/2, 7pi/4 d. pi/2, 7pi/6, 3pi/2, 11pi/6 I believe the answer is b but im not positive

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

its using the theta(0less than or eqal to theta less than or equal to 2 pi)

jimthompson5910 (jim_thompson5910):

cos(theta) - tan(theta)*cos(theta) = 0 cos(theta) - (sin(theta)/cos(theta))*cos(theta) = 0 cos(theta) - sin(theta) = 0 cos(theta) = sin(theta) cos^2(theta) = sin^2(theta) cos^2(theta) = 1-cos^2(theta) cos^2(theta)+cos^2(theta) = 1 2cos^2(theta) = 1 cos^2(theta) = 1/2 cos(theta) = sqrt(1/2) or cos(theta) = -sqrt(1/2) cos(theta) = sqrt(2)/2 or cos(theta) = -sqrt(2)/2 theta = arccos(sqrt(2)/2) or theta = arccos(-sqrt(2)/2) theta = pi/4 or theta = 5pi/4 ... use a unit circle here So you are correct.

OpenStudy (anonymous):

awesome thank you so much

jimthompson5910 (jim_thompson5910):

np